OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 25b2 - 15b + 9 16x2 - 56x + 49 9a2 - 20a - 25 25x2 - 40x - 16

OpenStudy (anonymous):

Do u understand or are you completely lost?

OpenStudy (anonymous):

completely lost

OpenStudy (anonymous):

can any one help me

OpenStudy (anonymous):

Answer: 16x2 − 56x + 49

OpenStudy (mathstudent55):

This is how you check it. A perfect square binomial is of the form: \(a^2 + 2ab + b^2 = (a + b)^2 \) or \(a^2 - 2ab + b^2 = (a - b)^2 \) Notice that in either case, for the trinomial the first term, a^2, is the square of a. The third term of the trinomial, b^2, is the square of b. The second term, 2ab, or -2ab is the product of 2, a, and b or -b.

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

welcomes :)

OpenStudy (mathstudent55):

The practical way of using the above information is like this. When you are given a trinomial to determine if it is a perfect square trinomial, follow these steps: 1. Is the first term the square of something? If so what? 2. Is the third term the square of something? If so what? 3. Is the second term 2 times the first term times the second term you found in steps 1 and 2 above? If all answers are yes, then it is a perfect square trinomial. If the answer to any of these questions is no, then it is not a perfect square trinomial.

OpenStudy (mathstudent55):

Look at choice A: 25b^2 - 15b + 9 1. 25b^2 is the square of 5b 2. 9 is the square of 3 and -3 3. Is 15b equal to 2 * 5b * 3 or 2 * 5b * (-3) ? 2 * 5b * 3 = 30b; 2 * 5b * (-3) = -30b The answer is no, so this is not a perfect square trinomial.